OpenStudy (narissa):

simplify the expression -4(x-3)+7(x+3)

OpenStudy (narissa):

@Austin1617

OpenStudy (narissa):

i have to show each step

pooja195 (pooja195):

Distribute: =(−4)(x)+(−4)(−3)+(7)(x)+(7)(3) =−4x+12+7x+21 Combine Like Terms: =−4x+12+7x+21 =(−4x+7x)+(12+21) Finish it off.

OpenStudy (narissa):

thank u

pooja195 (pooja195):

:)

OpenStudy (anonymous):

Wow @pooja195 was too fast

OpenStudy (narissa):

would it be 3x+33

OpenStudy (narissa):

@pooja195

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

but

pooja195 (pooja195):

@Austin1617 thats incorrect

OpenStudy (anonymous):

Its a expression so its 3x=33 I know @pooja195 !

OpenStudy (anonymous):

I said but @pooja195

OpenStudy (anonymous):

ok just nvm

OpenStudy (anonymous):

I'm sorry @pooja195 D: calm down

pooja195 (pooja195):

=(−4x+7x)+(12+21) -4+7=3 so 3x + <---the plus sign is here. NOT an EQUAL sign 12+21 = 33 :=) so therefore the answer is 3x+33 as @narissa stated.

OpenStudy (anonymous):

Just I know. I was going to explain narissa the same thing

OpenStudy (narissa):

im so confused thanks both of u

pooja195 (pooja195):

@narissa u were correct! lol

OpenStudy (anonymous):

Ok @narissa next question!